### v1

# Variational quantum state preparation for quantum-enhanced metrology in noisy systems

Abstract
We investigate optimized quantum state preparation for quantum metrology applications in noisy environments. We simulate a low-depth variational quantum circuit (VQC) composed of a sequence of global rotations and entangling operations applied to a chain of qubits that are subject to dephasing noise. The parameters controlling the VQC are numerically optimized to maximize the quantum Fisher information, which characterizes the ultimate metrological sensitivity of a quantum state, with respect to a global rotation. We find that regardless of the details of the entangling operation implemented in the VQC, the optimal quantum states can be broadly classified into a trio of qualitative regimes – cat-like, squeezed-like, and product states – associated with different dephasing rates. Our findings are relevant for designing optimal state-preparation strategies for next-generation quantum sensors exploiting entanglement, such as time and frequency standards and magnetometers, aimed at achieving state-of-the-art performance in the presence of noise and decoherence.
I Introduction
Quantum sensing holds great promise amongst emerging quantum technologies that exploit entanglement and quantum correlations as the closest to achieving a near-term, substantive quantum advantage over state-of-the-art classical devices. Prominent successes include LIGO, which has increased its detection volume with frequency-dependent squeezed light LIGO O4 Detector Collaboration et al. (2023), and the use of squeezed quantum states in searches for dark-matter candidates Jewell et al. (2023). Similarly, optical frequency metrology experiments are at the precipice of benefiting from quantum entanglement and correlations to push measurement precision beyond the limits of quantum projection noise Colombo et al. (2022); Robinson et al. (2024); Bothwell et al. (2022); Pedrozo-Peñafiel et al. (2020).
Progress towards the paradigm of quantum-enhanced sensing is being fueled through rapid improvements in technical capabilities, such as the realization of fully controllable intermediate scale quantum hardware featuring, e.g., 10s to even 1000s of high-quality qubits Ebadi et al. (2021); Manetsch et al. (2024). These systems exhibit long coherence times, an exquisite degree of coherent single-qubit control, and innovative methods for many-body control (i.e., entangling operations) via natural or engineered interactions Britton et al. (2012); Katz et al. (2023); Bornet et al. (2023a); Miller et al. (2024); Zhang et al. (2023).
A key challenge for current research is to establish robust metrological schemes that exhaust the potential of these hardware platforms in the presence of technical imperfections and decoherence. In the absence of these, optimal entangled quantum states for metrology are often exactly known, and can be obtained by analytically maximizing the quantum Fisher information $F_{Q}$, which bounds from below the achievable variance $(\Delta\theta)^{2}$ in an estimation parameter $\theta$ via the quantum Cramer-Rao bound, $(\Delta\theta)^{2}\geq 1/F_{Q}$ Braunstein and Caves (1994); Skotiniotis et al. (2015); Liu et al. (2019). In the context of sensing spin rotations, there exist a diversity of protocols, including non-equilibrium and quasi-adiabatic dynamics Lee (2006); Leibfried et al. (2005); Micheli et al. (2003); Nolan and Haine (2018), to prepare these ideal entangled states as well as optimize signal acquisition and readout to infer $\theta$ Leibfried et al. (2004); Haine (2018), and which can be tailored for specific experimental platforms. However, these optimal quantum states – corresponding to cat-like macroscopic superposition states – are catastrophically susceptible to decoherence Huelga et al. (1997); Pezzè et al. (2018), which is ever-present in real-world quantum hardware. It remains an open challenge to identify and prepare states that maximize quantum Fisher information in the presence of decoherence.
In this work, we perform a preliminary study of optimal state preparation for metrology in a noisy quantum environment. Our work is motivated by an effort to bring to bear existing resources and coherent control to maximize the potential of near-term quantum devices to demonstrate a state-of-the-art quantum advantage in sensing tasks, as opposed to proof-of-principle realizations. To make our task concrete, we numerically optimize a simple, low-depth variational quantum circuit (VQCs, see Fig. 1) for state preparation in an effort to maximize the resulting quantum Fisher information. Motivated by current physical platforms for quantum computation and simulation, we include a range of entangling gates in our VQC based on a variety of natural or engineered two-qubit interactions. We seek to maximize the QFI of the state generated by the VQC for a range of experimentally relevant noise models. We find that the optimal states can typically be grouped into a number of generic regimes – such as cat-like and squeezed-like quantum states – regardless of the nature of the entangling interactions, with some caveats associated with interaction range and system size. While our calculations are limited to relatively small systems (because the QFI for mixed quantum states is computationally expensive to obtain), we comment on implications for larger ensembles of qubits.
Our paper is structured as follows. In Sec. II we outline our basic VQC, including fundamental entangling operations and sources of decoherence. We present our results for the optimal QFI obtained with these VQCs in Sec. III, which includes a detailed analysis of the structure of the obtained optimal quantum states as well as a discussion of their dependence on system size and interaction range. We summarize our findings and present an outlook in Sec. IV.
II Variational quantum state preparation
[…]
As a basis for our investigation, we consider a low-depth VQC to prepare a quantum state of $N$ qubits to be used for sensing global qubit rotations. The applications for such states are diverse and include, e.g., frequency metrology and magnetometry. The VQC consists of two key pieces: (i) a set of global rotations occurring during the initial and final stages to prepare and reorient the qubits for sensing, and (ii) an intermediate stage where the qubits interact to create entanglement. In more detail (see Fig. 1), the VQC sequence begins by preparing an ensemble of $N$ qubits in a pure product state |$\ket{\psi_{0}}=\ket{\downarrow}^{\otimes N}$. The initial and final rotations are parameterized by $\hat{R}_{\theta_{j}}^{\varphi_{j}}=\prod_{k=1}^{N}e^{-i\theta_{j}\hat{\sigma}^{\varphi_{j}}_{k}/2}$ where $\hat{\sigma}_{k}^{\varphi_{j}}=\hat{\sigma}^{x}_{k}\cos{(\varphi_{j})}+\hat{\sigma}_{k}^{y}\sin{(\varphi_{j})}$ are Pauli matrices for the $k$th qubit in a rotated basis where $\theta_{j}$ and $\varphi_{j}$ define the rotation angle and axis for the initial ($j=1$) or final ($j=2$) rotation. The intermediate entangling step is generated by the gate:
$\hat{U}_{\theta_{\mathrm{I}}}=e^{-it_{\mathrm{I}}\hat{H}_{\mathrm{I}}(\chi)},$ (1)
where the Hamiltonian $\hat{H}_{\mathrm{I}}$ describes the interaction of the qubits with characteristic coupling strength $\chi$ (see below). We parameterize the “strength” of the entangling gate through the quantity $\theta_{\mathrm{I}}=\chi t_{\mathrm{I}}$ where $t_{\mathrm{I}}$ is the duration of the applied gate. Note that this preliminary description of the entangling gate as a unitary operation ignores the role of decoherence, but we will address this in the following subsections.
In total, the VQC is characterized by a set of five parameters $\mathbf{x}=(\theta_{1},\varphi_{1},\theta_{I},\theta_{2},\varphi_{2})$ and additionally through the choice of entangling Hamiltonian, $\hat{H}_{\mathrm{I}}$. We denote the output state of the VQC in terms of these parameters via the density matrix $\hat{\rho}(\mathbf{x})$.
II.1 Entangling gates
We study a variety of Hamiltonians $\hat{H}_{\mathrm{I}}$ for the entangling gate, motivated by both conceptual and practical considerations. Firstly, to benchmark our findings we explore one-axis (OAT) and two-axis (TAT) twisting Hamiltonians featuring infinite-range pairwise interactions between qubits Kitagawa and Ueda (1993). The generation of entangled states for metrology using the unitary dynamics generated by these Hamiltonians, such as spin-squeezed and Greenberger-Horne-Zeilinger (GHZ) states, is well established Kitagawa and Ueda (1993); Tóth and Apellaniz (2014); Pezzè et al. (2018). In particular, OAT dynamics have been realized and exploited in a diversity of experimental platforms including trapped ions Bohnet et al. (2016), Bose-Einstein condensates Muessel et al. (2014) and cavity-QED Leroux et al. (2010); Hosten et al. (2016); Norcia et al. (2018); Braverman et al. (2019), and there has been early work proposing He et al. (2019) and even demonstrating Miller et al. (2024); Luo et al. (2024) TAT interactions. In addition, motivated by current efforts to generate entangled states in quantum simulation platforms featuring finite-range pairwise couplings between qubits Perlin et al. (2020); Young et al. (2023); Eckner et al. (2023); Franke et al. (2023); Bornet et al. (2023b), we also study Ising, XX, and f-TAT (finite-range TAT) spin chain models with power law interactions. The Hamiltonians for these models are
$\hat{H}_{\mathrm{Ising}}$ $=\sum_{i\neq j}{\chi_{ij}\hat{\sigma}_{i}^{z}\hat{\sigma}_{j}^{z}},$ (2)
$\hat{H}_{\mathrm{XX}}$ $=\sum_{i\neq j}{\chi_{ij}(\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}+\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{y}}),$ (3)
$\hat{H}_{\text{f-TAT}}$ $=\sum_{i\neq j}{\chi_{ij}(\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{y}+\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{x})}.$ (4)
Here $i$ and $j$ index individual qubits at positions $r_{i}$ and $r_{j}$, and the couplings $\chi_{ij}=\chi/\mathcal{N}_{\alpha}|r_{i}-r_{j}|^{\alpha}$, where $\chi$ sets the strength of interactions and $\mathcal{N}_{\alpha}=(N-1)^{-1}\sum_{i\neq j}|r_{i}-r_{j}|^{-\alpha}$ is a normalization factor to ensure that the Hamiltonian is extensive Uhlenbeck et al. (1963); Defenu (2021) and simplify comparison of the models as a function of decoherence strength (see below). Throughout this manuscript we set $\hbar=1$ and, for simplicity, we focus on qubits in a one-dimensional geometry that can be realized using, e.g., arrays of neutral atoms or trapped ions.
We will restrict our focus to interactions featuring power-law exponents $\alpha=0,...,6$, motivated by the range featured in prominent quantum science platforms such as Rydberg atoms ($\alpha=3,6$), trapped ions ($\alpha\in[0,3]$), molecules ($\alpha=3$) and light-matter systems ($\alpha=0$). The limiting case of $\alpha=0$ corresponds to all-to-all interactions, for which the Ising and XX models reduce (up to a constant of motion) to the OAT Hamiltonian $\propto\hat{S}_{z}^{2}$, while f-TAT resolves to the TAT Hamiltonian $\propto(\hat{S}_{x}\hat{S}_{y}+\hat{S}_{y}\hat{S}_{x})$.
II.2 Dephasing noise
Motivated by the implementation of the above VQC in real-world hardware, we incorporate into our description decoherence due to, e.g., stray background fields and spurious couplings to the environment. However, in our treatment we will make the simplifying assumption that decoherence only occurs during the application of the entangling gate, with the dynamics of the initial and final rotations taken to be sufficiently rapid that they may be approximated as unitary processes.
The inclusion of decoherence and dissipation means that the unitary representation of the entangling gate given by Eq. (1) is inappropriate. Instead, we model this stage of the VQC using a Linblad master equation,
$\frac{d\hat{\rho}}{dt}=-i[\hat{H}_{\mathrm{I}},\hat{\rho}]+\sum_{\nu,j}\gamma_{\nu}\left(\hat{L}^{\nu}_{j}\hat{\rho}\hat{L}^{\nu\dagger}_{j}-\frac{1}{2}\{\hat{L}^{\nu\dagger}_{j}\hat{L}^{\nu}_{j},\hat{\rho}\}\right).$ (5)
where $\hat{H}_{\mathrm{I}}$ is the chosen entangling Hamiltonian [Eqs. (2)-(4)], $\nu\in\set{x,\;y,\;z}$ indexes an axis for dephasing at rate $\gamma_{\nu}$, and $j\in\set{1,2,\cdots,N}$ indexes a qubit addressed by the jump operator $\hat{L}^{\nu}_{j}=\hat{\sigma}^{\nu}_{k}/2$. In this work, we will assume that the dephasing is isotropic, i.e., acts with equal rates $\gamma_{x}=\gamma_{y}=\gamma_{z}=\gamma$. This assumption avoids bias in our results due to, e.g., robustness of the entangling dynamics for specific representations of the Hamiltonians in Eqs. (2)-(4Britton et al. (2012); Jurcevic et al. (2014). In future work Zuñiga Castro et al. (ming), we will consider the generic case of anisotropic dephasing (unequal $\gamma_{x},\gamma_{y},\gamma_{z}$) as well as decoherence due to spontaneous emission and absorption.
II.3 Optimization of the quantum Fisher information
We seek to identify optimal state preparation protocols, within the constraints of the VQC defined above, for metrology of global spin rotations. The quantity capturing the suitability of a generic quantum state $\hat{\rho}$ for this task is the quantum Fisher information Pang and Brun (2014)
$F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{\lambda_{k}+\lambda_{l}}|\braket{k}{\hat{G}}{l}|^{2}.$ (6)
Here |$\ket{k}$ and $\lambda_{k}$ are, respectively, the eigenvectors and corresponding eigenvalues of the density matrix prepared by the VQC, i.e., $\hat{\rho}=\sum_{k}\lambda_{k}\ket{k}\!\bra{k}$ |, and $\hat{G}$ is the generator of a unitary transformation $\hat{U}_{\phi}=e^{-i\phi\hat{G}}$ encoding a classical parameter $\phi$. The QFI characterizes the uncertainty with which the classical parameter can be estimated through the quantum Cramer-Rao bound (QCRB), $(\Delta\phi)^{2}\geq 1/F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$. Without loss of generality, throughout this work we set $\hat{G}=\hat{S}_{z}=\sum_{j}\hat{\sigma}^{z}_{j}/2$.
[…]
|$\ket{\mathrm{GHZ},\Phi}=(\ket{\uparrow}^{\otimes N}+e^{i\Phi}\ket{\downarrow}^{\otimes N})/\sqrt{2},$ (7)
with arbitrary phase $\Phi$. Strategies to dynamically prepare this state are known and have even been demonstrated for small systems Leibfried et al. (2005), although their feasibility for scalable quantum-enhanced sensing is severely limited by the catastrophic sensitivity of the cat state to decoherence Huelga et al. (1997); Foss-Feig et al. (2013).
In the following section, we instead investigate optimal preparation strategies, using the VQC discussed previously, that take into account the effects of decoherence. For a fixed decoherence strength $\gamma$, we obtain the state $\hat{\rho}(\mathbf{x})$ generated by a given VQC with parameters $\mathbf{x}$. We numerically integrate the master equation (5) for a given choice of entangling Hamiltonian $\hat{H}_{\mathrm{I}}$ and associated global rotations. The optimal preparation strategies defined by parameters $\mathbf{x}_{\mathrm{opt}}$ are obtained by numerically optimizing over $\mathbf{x}$ to maximize the QFI of Eq. (6). Throughout this work we use derivative-free algorithms such as the Nelder-Mead method in SciPy Virtanen et al. (2020) to obtain $\mathbf{x}_{\mathrm{opt}}$. The code package used to simulate the VQC is available at Ref. Colussi et al. (2024).
III Results
In this section, we present results of our numerical optimization for the range of entangling Hamiltonians defined in Eqs. (2)-(4) and for extreme cases of $\alpha=0$ and $6$. A key result of our investigation is that we observe that the optimal states prepared by the VQC can be typically delineated into three regimes: (i) perturbative, characterized by weak decoherence such that the optimal preparation strategy generates macroscopic superposition states that approach the cat state that is known to achieve sensitivity at the HL in the absence of noise, (ii) squeezed-like, characterized by states with non-zero polarization (i.e., collective spin projection) and anisotropic projection noise, and (iii) uncorrelated, corresponding to a separable coherent spin state of the qubits with sensitivity near the SQL. This classification is based on analysis of the optimal QFI and abrupt transitions in the associated VQC parameters, but is further supported through the examination of a number of auxiliary quantities, including state overlap, distribution functions and spin squeezing, which provide additional information about the structure of the prepared quantum states.
Figure 2: The normalized QFI $F_{Q}/N^{2}$ and optimal parameters $\mathbf{x}_{\mathrm{opt}}$ as functions of isotropic decoherence strength, $\gamma$ for (a) OAT and TAT, and (b) Ising, XX and f-TAT models. In all panels we use $N=8$ and the interaction range is set to (a) $\alpha=0$ and (b) $\alpha=6$. For all cases we exclude the initial rotation angle $\theta_{1}$, as it is found to identically be either $\pi/2$ (OAT, Ising and XX) or 00 (TAT and f-TAT) regardless for non-trivial values of $\gamma<\gamma^{(2)}$ (see main text). The values of $\gamma^{(1)}$ and $\gamma^{(2)}$ for each model are indicated by vertical lines (blue dashed for OAT and Ising, red dot-dashed for TAT and f-TAT and solid green for XX) to guide the reader between the main regimes. Note that $\gamma^{(1)}$ is absent for the Ising data, while the values of $\gamma^{(1)}$ for XX and f-TAT are indistinguishable.
III.1 One- and two-axis twisting
We begin by presenting results for OAT and TAT with infinite-range interactions ($\alpha=0$), as the dynamics generated by these Hamiltonians is well understood in the absence of noise Kitagawa and Ueda (1993); Kajtoch and Witkowska (2015) and they can provide a template for our later discussion of finite-range interactions. The optimal QFI and parameters obtained for OAT and TAT are plotted in panel (a) of Fig. 2 as a function of the isotropic dephasing strength $\gamma$ and with fixed $N=8$. For both cases, the QFI begins near the maximal value $F_{Q}/N^{2}\approx 1$ (HL) for $\gamma\ll 1$ (in fact, it is essentially saturated for OAT) and then monotonically decreases until settling at the SQL $F_{Q}/N^{2}=1/N$ for $\gamma\gg 1$.
We observe that the QFI appears to be continuous as function of $\gamma$ for both OAT and TAT, although there is a clear kink in the QFI at $\gamma\approx 0.25$ for OAT. On the other hand, the associated VQC parameters $\mathbf{x}_{\mathrm{opt}}$ show more abrupt behaviour, including discontinuities, that we use to qualitatively classify different regimes of quantum states. In particular, we use the strength of the entangling gate $\theta_{\mathrm{I}}$ to define two characteristic values of $\gamma$. For vanishingly small $\gamma$ we observe that $\theta_{I}$ takes the largest values for both OAT and TAT, but it slowly decreases with increasing $\gamma$. We identify $\gamma^{(1)}$ as the dephasing strength at which $\theta_{I}$ exhibits a discontinuity and abruptly drops in value, leading to $\gamma^{(1)}\approx 0.25$ and $0.02$ for OAT and TAT, respectively. Further decrease in $\theta_{I}$ is then observed for $\gamma>\gamma^{(1)}$ until it vanishes, i.e., the optimal VQC does not feature any application of the entangling gate. We define the point at which $\theta_{I}=0$ as $\gamma^{(2)}$, and find $\gamma^{(2)}\approx 9.44$ and $25.85$ for OAT and TAT, respectively. Additionally, we observe that similar discontinuities or abrupt behaviour are present in other VQC parameters, e.g., rotation angle $\theta_{2}$ and the axis difference $|\varphi_{1}-\varphi_{2}|/\pi$. Note that we do not show data for the initial rotation angle $\theta_{1}$, as we find that the optimal choice is $\theta_{1}=\pi/2$ for OAT and $\theta_{1}=0$ for TAT for $\gamma<\gamma^{(2)}$ (i.e., “non-trivial” values of $\gamma$).
Insight into the three distinct parameter regimes carved out by the definition of $\gamma^{(1)}$ and $\gamma^{(2)}$ is found by drilling further into the OAT results, which can be contrasted against the well-understood unitary limit of $\gamma=0$. Firstly, when $\gamma<\gamma^{(1)}$ we observe that the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ [see Fig. 2 panel (a)] are virtually identical to those that would produce a GHZ state [see Eq. (7)] in the absence of decoherence Foss-Feig et al. (2013). In particular, our optimal VQC sequence corresponds to a first rotation of $\theta_{1}=\pi/2$ with arbitrary value of $\varphi_{1}$, which orients the initially prepared product state on the collective Bloch sphere equator. The rotation is followed by an entangling gate with strength $\theta_{\mathrm{I}}\approx\pi/2$, which is consistent with the value required to generate a GHZ state using twisting dynamics Foss-Feig et al. (2013). At this point, the generated GHZ-like state is characterized by a collective superposition of qubits oppositely polarized along an axis lying in the equatorial plane (specifically, along an axis perpendicular to that defined by the first rotation about $\varphi_{1}$). Thus, the final rotation by an angle $\theta_{2}=\pi/2$ about the axis $\varphi_{2}\approx\varphi_{1}$ reorients the GHZ-like state along the $z$ axis [similar to Eq. (7)]. In the absence of noise, this state features the largest possible fluctuations in the observable $\hat{S}_{z}$ and thus maximizes the QFI Tóth and Apellaniz (2014). As a result, we dub $\gamma<\gamma^{(1)}$ as the perturbative regime, as our results suggest that decoherence is sufficiently weak so that the optimal state preparation strategy is simply to closely (i.e., perturbatively) follow the approach established in the absence of decoherence.
On the other hand, in the regime where $\gamma^{(1)}<\gamma<\gamma^{(2)}$, $\theta_{\mathrm{I}}$ takes on relatively smaller values that can be associated with the paradigmatic generation of spin-squeezed states Kitagawa and Ueda (1993). Particular to one-axis twisting, the unitary generation of spin-squeezed stated occurs via “shearing” of the quantum fluctuations of an initial coherent spin state polarized along the equator of the collective Bloch sphere. The resulting squeezing of the projection noise is thus in a spin quadrature that is rotationally offset from the equatorial plane of the Bloch sphere, by an amount that depends on the strength/duration of the twisting interaction ($\theta_{I}$), and thus requires a final rotation about an axis aligned along the collective polarization of the state. We see signatures of these features in our dynamics with decoherence, including that the final rotation axis is perpendicular to the initial rotation (and thus parallel to the polarization of the state post first rotation), $|\varphi_{2}-\varphi_{1}|\approx\pi/2$, whilst the angle of the rotation $\theta_{2}$ varies with the duration of the entangling interaction characterized by $\theta_{\mathrm{I}}$ Kitagawa and Ueda (1993). Thus, since our results suggest that the states within this regime may be closely associated with spin-squeezing we designate $\gamma^{(1)}<\gamma<\gamma^{(2)}$ as the squeezed-like regime.
Finally, we denote $\gamma>\gamma^{(2)}$ as the uncorrelated regime. In this case, as the entangling gate is never applied ($\theta_{I}=0$), the VQC reduces to a pair of global rotations that together simply reorient the initial pure, uncorrelated coherent spin state onto some axis in the $x$-$y$ plane to maximize the QFI at the SQL, $F_{Q}/N^{2}=1/N$. This regime arises for overwhelmingly large decoherence, such that any application of the entangling gate would in fact lead to a QFI below the SQL due to an increase in total projection noise, i.e., the interactions cannot generate metrologically useful correlations faster than decoherence degrades the quantum state [see the bottom subplot of Fig. 3].
Figure 3: Diagnostic quantities calculated for optimally prepared states as functions of dephasing strength $\gamma$ for (a) OAT and (b) TAT. All data is for $N=8$. (Top subplots) Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs) and the normalized QFI ($F_{Q}/N^{2}$, circles for OAT and triangles for TAT). (Lower subplots) The overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S(S+1)$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}$ and $\gamma^{(2)}$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For the GHZ-like states we show two perspectives of the same quantum state, rotated by $\pi/2$ about indicated axes.
The behaviour of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ for TAT (also shown in Fig. 2) follow similar trends and is suggestive of a similar classification into three state preparation regimes. However, in this case the increased complexity of the unitary dynamics for TAT Kajtoch and Witkowska (2015) makes quantitative analysis of the parameters more difficult. Thus, in the following we build on the intuition developed from OAT and perform a more systematic investigation of the structure of the prepared quantum states using a variety of diagnostic quantities.
III.1.1 Characterization of GHZ-like states
We first focus on the perturbative regime defined by $\gamma<\gamma^{(1)}$. Motivated by the observation that the QFI in this regime is near the HL ($F_{Q}/N^{2}=1$) and informed by the close proximity of the parameters $\mathbf{x}_{\mathrm{opt}}$ for OAT to those that are known to generate GHZ states [see Eq. (7)] for $\gamma=0$, we calculate the overlap between the optimally prepared state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ and a GHZ state. Specifically, we compute the maximum fidelity,
$\mathcal{F}_{\mathrm{GHZ}}=\max_{\Phi}~{}|\braket{\mathrm{GHZ},\Phi}{\hat{\rho}(\mathbf{x}_{\mathrm{opt}})}{\mathrm{GHZ},\Phi}|^{2}.$ (8)
The results for this quantity as a function of $\gamma$ are shown in the middle panels of Fig. 3(a) and (b). We observe appreciable fidelities $1/2<\mathcal{F}_{\mathrm{GHZ}}\leq 1$ in the regime $\gamma<\gamma^{(1)}$ for both OAT [panel (a)] and TAT [panel (b)]. The values for the fidelity are indicative of strong overlap with the GHZ state and the presence of $N$-qubit entanglement Leibfried et al. (2005). For both OAT and TAT the fidelity remains appreciable when $\gamma>\gamma^{(1)}$, although it steadily decreases Huelga et al. (1997). In particular, the fidelity appears to be more robust to increasing decoherence for TAT. However, we qualify these observations by noting that the system size used in our calculations is quite small.
The fidelity is qualitatively supported by visualization of the quantum states in this regime using an averaged Husimi probability distribution (see Appendix A), shown below the main panels of Fig. 2. These distributions clearly demonstrate that the quantum state is composed of a superposition of two components localized at the north and south pole of the Bloch sphere, consistent with a GHZ-like state.
III.1.2 Characterization of squeezed-like states
In the squeezed-like regime found for larger decoherence, $\gamma^{(1)}<\gamma<\gamma^{(2)}$, we expect the prepared states to have a non-zero collective polarization and anisotropic quantum fluctuations in the spin quadratures defined perpendicular to the orientation of the state. Mathematically, this structure can be captured by the Wineland squeezing parameter Wineland et al. (1992),
$\xi^{2}_{s}=N\min_{\mathbf{n}_{\perp}}\frac{\langle(\Delta\hat{S}_{\mathbf{n}_{\perp}})^{2}\rangle}{|\langle\hat{\mathbf{S}}\rangle|^{2}},$ (9)
where $\langle\hat{\mathbf{S}}\rangle=(\langle\hat{S}_{x}\rangle,\langle\>\hat{S}_{y}\rangle,\>\langle\hat{S}_{z}\rangle)$ is the collective Bloch vector of the state and $\mathbf{n}_{\perp}$ denotes a unit vector lying in a plane orthogonal to $\langle\hat{\mathbf{S}}\rangle$. The squeezing parameter is obtained by minimizing the fluctuations $\langle(\Delta\hat{S}_{\mathbf{n}_{\perp}})^{2}\rangle$ with respect to the orientation of $\mathbf{n}_{\perp}$. A value $\xi^{2}_{s}<1$ reflects that the quantum fluctuations along one spin quadrature are squeezed below the projection noise limit such that the state can enable sensing of spin rotations below the SQL using a simple Ramsey sequence Wineland et al. (1992). More generally, squeezing sets a lower bound on the QFI (and thus metrological utility of the quantum state) through the inequality $F_{Q}/N^{2}\geq(N\xi^{2}_{s})^{-1}$.
We plot the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ as a function of $\gamma$ in the upper panels of Fig. 3(a) and (b). We observe that $(N\xi^{2}_{s})^{-1}=0$ in the perturbative regime of $\gamma<\gamma^{(1)}$ as the polarization of the cat state vanishes, i.e., $\langle\hat{\mathbf{S}}\rangle=0$. On the other hand, for $\gamma>\gamma^{(1)}$ we observe $(N\xi^{2}_{s})^{-1}>0$. Initially, as $\gamma$ is increased past this transition, the inverse squeezing parameter does not breach the SQL, i.e., $(N\xi^{2}_{s})^{-1}<1/N$, indicating that while the QFI is large ($F_{Q}/N^{2}>1/N$) and corresponds to a quantum-enhanced sensitivity, the state does not exhibit simple squeezing that can be exploited using Ramsey interferometry with mean values of global spin measurements. We attribute this to oversqueezing, which has been studied in the context of OAT dynamics Strobel et al. (2014); Bohnet et al. (2016). To demonstrate this, we show example Husimi distributions for representative quantum states in this regime below panels Fig. 3(a) and (b). These clearly illustrate that the quantum state features anisotropic quantum noise. However, the distribution wraps appreciably around the Bloch sphere, leading to a shrinking of the collective Bloch vector $\langle\hat{\mathbf{S}}\rangle$ and the failure of the squeezing parameter to capture the metrological sensitivity of the state. As $\gamma$ further increases, the squeezing parameter breaks through the SQL and eventually matches the QFI, $(N\xi^{2}_{s})^{-1}=F_{Q}/N^{2}$. In this latter regime, the metrological utility of the state is thus entirely captured by spin squeezing.
III.1.3 Additional diagnostic quantities
In addition to the fidelity to the GHZ state and the squeezing parameter, it is insightful to consider additional quantities. First, we compute the collectivity of the prepared states [see Fig. 3], $\langle\hat{S}^{2}\rangle$ which should be compared against the maximal value of $S(S+1)$ with $S=N/2$ for a collective state. Both OAT and TAT unitary dynamics preserve the collectivity as they feature infinite-range interactions, so that any deviation from the maximum value of $\langle\hat{S}^{2}\rangle$ should be attributable to the single-qubit dephasing noise. The collectivity is most useful examined alongside a second quantity, the classical Fisher information (CFI), which can be obtained as,
$F_{c}(\theta)=\lim_{\theta\to 0}\frac{4}{\theta^{2}}\sum_{m_{z}}\left(\sqrt{P(m_{z};\theta)}-\sqrt{P(m_{z};\theta=0)}\right)^{2}$ (10)
where $P(m_{z})$ is the distribution function for the collective spin projection $m_{z}$ along $\hat{z}$, and $F_{c}$ is bounded from above by the QFI, $F_{c}\leq F_{Q}$. In the case of unitary OAT and TAT dynamics, the CFI always saturates the QFI when optimized over all possible collective measurements. However, we observe $F_{c}$ is less than the QFI across parts of both the perturbative and squeezed-like regimes, particularly when the collectivity strays appreciably from $S(S+1)$. This suggests that more sophisticated measurements, such as local (site-resolved) spin observables are necessary to exhaust the metrological potential of the states we preserve, and in particularly highlights that states in the “squeezed” regime – a name that is motivated by a simple analysis of collective spin fluctuations – require more nuanced treatment in general.
III.2 Ising, XX, f-TAT
Figure 4: Diagnostic quantities calculated for optimally prepared states as functions of dephasing strength $\gamma$ for (a) Ising, (b) XX and (c) f-TAT models. All data is for $N=8$ and $\alpha=6$. (Top subplots) Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs) and the normalized QFI ($F_{Q}/N^{2}$, circles, diamonds and triangles, respectively). (Lower subplots) The overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S(S+1)$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}$ [absent for (a)] and $\gamma^{(2)}$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For the GHZ-like states we show two perspectives of the same quantum state, rotated by $\pi/2$ about indicated axes.
Equipped with an understanding of the results for the infinite-ranged OAT and TAT entangling Hamiltonians, we now generalize to the finite-ranged models given in Eqs. (2)-(4). For simplicity, we focus on the specific coupling exponent $\alpha=6$ in the following, and will present additional results investigating the role of interaction range in a more detailed follow-up work Zuñiga Castro et al. (ming).
In Fig. 2(b), we plot the maximal QFI and optimal parameters $\mathbf{x}_{\mathrm{opt}}$ as a function of dephasing strength $\gamma$ with $N=8$ qubits. Overall, we observe a similar delineation of the QFI and $\mathbf{x}_{\mathrm{opt}}$ into different regimes set by the dephasing strength, with some important caveats. Firstly, for the XX and f-TAT models, we define an analogous perturbative regime for $\gamma<\gamma^{(1)}$ (see vertical lines in panels), where the transition $\gamma^{(1)}$ is again defined by an abrupt drop in the interaction time characterized by $\theta_{I}$ and similarly sharp changes in the rotation angles/axes. Similarly, we define a squeezed-like regime for $\gamma^{(1)}<\gamma<\gamma^{(2)}$, with the upper bound on the dephasing strength $\gamma^{(2)}$ set via the uncorrelated regime that emerges when $\theta_{I}$ finally reaches zero. On the other hand, the Ising model presents distinct behaviour. In particular, we observe an absence of the perturbative regime, as the VQC parameters change smoothly throughout a large squeezed-like regime $\gamma<\gamma^{(2)}$, until $\theta_{I}$ vanishes at the boundary of the uncorrelated regime that exists for $\gamma\geq\gamma^{(2)}$.
While our preliminary identification of the perturbative, squeezed-like and uncorrelated regimes is supported by an analysis of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$, which show some qualitative similarities to OAT (Ising and XX) and TAT (f-TAT), we provide stronger evidence through quantities such as the overlap with a GHZ state and the squeezing parameter, which are plotted as a function of $\gamma$ in Fig. 4.
The perturbative regime for the XX and f-TAT models [see panels (b) and (c)] is supported by an appreciable overlap $\mathcal{F}_{\mathrm{GHZ}}$ with the GHZ state. However, unlike the infinite-range models, the overlap dips below a value of $1/2$ before transitioning to the squeezed-like regime. This shows that, while in the limit $\gamma\to 0$ the short-range interactions can generate quasi-collective entangled states (i.e., $\mathcal{F}_{\mathrm{GHZ}}>1/2$) Perlin et al. (2020), this eventually degrades as $\gamma$ is increased. Further to this, we highlight that the characteristic collectivity $\langle\hat{S}^{2}\rangle$ of the states generated in this regime is a substantial fraction of the maximum value $S(S+1)$. In contrast, the Ising model [see panel (a)] never features a meaningfully large overlap with the GHZ state, even for $\gamma\to 0$, although we note the value of the collectivity $\langle\hat{S}^{2}\rangle$ is not too different from the states prepared by XX and f-TAT.
The squeezed-like regime between $\gamma^{(1)}<\gamma\leq\gamma^{(2)}$ is more complex for the short-ranged models. While the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ takes non-zero values for $\gamma>\gamma^{(1)}$, reflecting the transition from the GHZ state to one with non-zero collective polarization of the qubits, in contrast to the results for the infinite-ranged models we observe that the squeezing parameter never breaches the SQL (i.e., $(N\xi^{2}_{s})^{-1}<1/N$). This reflects that while these states may exhibit anistropic fluctuations/correlations analogous to a spin-squeezed state(see averaged Husimi distributions in each panel), the finite-range interactions lead to states that are far from collective. Notably, the collectivity parameter $\langle\hat{S}^{2}\rangle$ is far smaller in the squeezed-like regime for these models than previous observations for the infinite-ranged models in Fig. 3. This complexity of the squeezed-like regime is further supported by computation of the CFI. For the Ising model, there is a small regime where the CFI breaches the SQL, but for the XX and f-TAT models the CFI is effectively always at or just worse than the SQL. Given the CFI that we compute is constrained to collective observables, this suggests that harnessing the QFI of states generated by finite-ranged interactions will require measurement of site-resolved observables or correlations.
Figure 5: Dephasing strengths $\gamma^{(j)}$ ($j=1,2$) as a function of $N$ for (a) OAT and TAT ($\alpha=0$), and (b) Ising, XX and f-TAT ($\alpha=6$). Solid and dashed lines are used to guide the eye and distinguish $\gamma^{(1)}$ and $\gamma^{(2)}$, respectively.
III.3 Dependence on qubit number
Our results so far have focused on the specific case of $N=8$. However, it is important to examine the dependence of our analysis on system size, particularly to generate insight and develop general strategies for larger systems that are not accessible with exact numerical calculations.
We conduct additional calculations for system sizes $N=2,4,6,8$ and perform a similar analysis to that described in the previous sections. In particular, we still observe a general delineation of the optimal preparation strategy into the three regimes: perturbative, squeezed-like and uncorrelated. We extract the transition values $\gamma^{(1)}$ and $\gamma^{(2)}$ for each system size and plot the results in Fig. 5 (the solid and dashed lines are used to distinguish the markers for $\gamma^{(1)}$ and $\gamma^{(2)}$, respectively).
Panel (a) of Fig. 5 shows the results for OAT and TAT. We observe that for both entangling Hamiltonians, the scope of the perturbative regime tends to decrease with system size, i.e., $\gamma^{(1)}$ trends to smaller values. This is unsurprising given that GHZ-like states are known to be highly fragile and difficult to dynamically prepare in the presence of decoherence, and this fragility worsens with system size Foss-Feig et al. (2013); Huelga et al. (1997). We note that while $\gamma^{(1)}$ appears to monotonically decrease with system size for OAT, TAT exhibits more sophisticated behaviour. Specifically, TAT depicts a slight increase in $\gamma^{(1)}$ from $N=2$ to $4$ before dropping away very quickly relative to OAT. We attribute this to the fact that TAT does not prepare explicit GHZ states (in fact the characteristic fidelity $\mathcal{F}_{\mathrm{GHZ}}$ deep in the perturbative regime changes noticeably as $N$ is varied) and we are probing particularly small system sizes. In future work, it would be interesting to exploit the permutational symmetry of the TAT dynamics using efficient numerical methods Baragiola et al. (2010) to study the non-unitary dynamics for much larger systems $N\gtrsim 100$, in an effort to better understand the complexity of the perturbative regime. In contrast to the shrinking perturbative regime, the value of $\gamma^{(2)}$ increases similarly with system size for OAT and TAT. Thus for large $N$ we expect that, for this VQC, preparing squeezed-like states will become the clear optimal strategy for sensing spin rotations in the presence of modest decoherence.
Panel (b) of Fig. 5 shows analogous results for the Hamiltonians in Eqs. (2)-(4) with a choice of $\alpha=6$. Intriguingly, for the XX and f-TAT Hamiltonians we observe that $\gamma^{(1)}$ settles to a fixed value $\gamma^{(1)}\approx 0.46$ independent of system size for $N\geq 4$ (for the Ising Hamiltonian no perturbative regime exists so we do not plot $\gamma^{(1)}$). Note that this does not imply that finite-range interactions are more robust to decoherence, as the QFI of the states in this persistent perturbative regime can be much smaller than for infinite-range OAT and TAT (see Fig. 2). On the other hand, the increase of $\gamma^{(2)}$ with system size for each of the Ising, XX and f-TAT Hamiltonians closely follows the results for OAT and TAT.
IV Conclusion
With a focus on the preparation of states with optimal QFIs within a framework of noisy environments, our results confirm the existence of quantum states with QFIs above the SQL relevant to many fields, e.g., quantum metrology and quantum information science. We inform on the nature of the states generated with a set of physically relevant collective and finite-range Hamiltonians for a given decoherence $\gamma$ and classify them into three regimes closely associated with their QFIs and optimal preparation strategies. While we find that the generation of states with QFIs scaling near the HL appears to be an increasingly ambitious task with increasing system size (on account of rising sensitivities to deleterious noise), we emphasize that the generation of states with QFIs above the SQL scales favorably with $N$. Furthermore, our results indicate promise for the generation of states with QFIs practical toward quantum enhanced sensing and communication tasks with physical systems featuring limited interaction ranges.
There are a number of interesting future directions that can build on the analysis presented in this work. First, we can expand the circuit ansatz in Fig. 1. For example, we can increase the depth of our VQC by layering alternating rotations and entangling stages. This will, in principle, increase the complexity and range of potential states that we can prepare, potentially modifying our current classification into perturbed, squeezed-like, and uncorrelated states. The VQC ansatz can also be expanded to fully exploit the programmability of quantum sensing hardware by including different site-resolved qubit rotations and two-qubit gates, as opposed to global rotations and entangling operations.
Second, in addition to decoherence introduced during the entangling stage of the VQC, we can incorporate decoherence introduced by errors in VQC parameters such as imperfect initial and final rotations (i.e., shot-to-shot fluctuations in rotation axis or angle). In contrast to the current VQC that we study, this could benefit theoretical understanding by introducing some degree of control over the level of mixedness of the prepared states that is independent of the VQC parameters (whereas in the current protocol the total amount of decoherence is correlated with the duration of the entangling stage).
Third, our investigation of the QFI of mixed quantum states can be broadened by considering isolated quantum systems that feature multiple degrees of freedom that can be entangled and coherently controlled but are not necessarily mutually measurable. This leads to a situation where the metrological potential of a prepared quantum state is best captured by the QFI of the reduced density matrix describing the observable degree(s) of freedom, which will be mixed if the system features entanglement Lewis-Swan et al. (2024).
Finally, we are investigating the development of more sophisticated, specialized numerical optimization routines for optimizing the QFI for a given VQC or state preparation protocol. Whereas general-purpose optimization methods, such as the Nelder-Mead algorithm used to generate the results in this manuscript, are designed to receive only scalar outputs, the QFI form in Eq. (6) has considerable structure that can be exploited. Instead of receiving only the scalar value $F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$, we are developing optimization methods that can receive and work with the density matrix $\hat{\rho}(\mathbf{x})$ and approximate (finite-differences) or exact (algorithmic-differentiated) derivatives for entries of $\hat{\rho}(\mathbf{x})$ with respect to the parameters $\mathbf{x}$. It is not straightforward to pass this information through an eigendecompostion, which is needed in Eq. (6). However, by building on past work by some of us Larson et al. (2021); Larson and Menickelly (2024) that optimizes composite objectives by using objective knowledge form (such as the form of Eq. (6)) in a QFI-only numerical optimization routine, we hope to find higher-quality solutions in fewer queries/simulations of the system. This is a particularly important task, as simulation of open quantum systems involving many qubits and subsequent computation of the QFI for a given mixed quantum state is computationally expensive.
V Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, Exploratory Research for Extreme-Scale Science program.
Appendix A Husimi probability distributions
Here we define the Husimi probability distributions used to visualize spin states in Fig. 3. The basic idea is to define a Husimi probability distribution within each manifold of fixed spin length $S$, and take an appropriately weighted average of these distributions. Note that for ease of language, in this appendix we treat the symbol $S$ as a variable to denote spin length, whereas the main text defines $S=N/2$ for $N$ spins.
A.1 A simplified case
Consider first the simplified case in which $N$ spins occupy the permutationally symmetric manifold with spin length $S=N/2$. In this case, for a density matrix $\hat{\rho}$ we can define the Husimi probability distribution
$\rho(\bm{v})=w_{S}\braket{\bm{v}}{\hat{\rho}}{\bm{v}}$ (11)
|$\ket{\bm{v}}=\hat{R}(\bm{v})\ket{\uparrow}^{\otimes N},$ $\hat{R}(\bm{v})=e^{-i\phi\hat{S}_{z}}e^{-i\theta\hat{S}_{y}},$ (12)
and the normalization factor $w_{S}$ is determined by requiring that $\rho(\bm{v})$ is a normalized probability distribution, with
$\int\text{d}\bm{v}\,\rho(\bm{v})=\int_{0}^{\pi}\text{d}\theta\,\sin\theta\int_{0}^{2\pi}\text{d}\phi\,\rho(\theta,\phi)=1,$ (13)
which implies that [^1] The normalization factor $w_{S}$ is most easily determined by considering a uniform mixture of all $N+1$ permutationally symmetric states, for which $\braket{\bm{v}}{\hat{\rho}}{\bm{v}}=1/(N+1)=1/(2S+1)$.
$w_{S}=\frac{2S+1}{4\pi}.$ (14)
When the state $\hat{\rho}$ is permutationally symmetric, the probability distribution $\rho$ is a faithful representation of $\hat{\rho}$ in the sense that $\hat{\rho}$ is uniquely determined by $\rho$. Moreover, in this case the values $\rho(\bm{v})$ at $(2S+1)^{2}$ points $\bm{v}$ suffice to reconstruct $\hat{\rho}$ Perlin et al. (2021).
A.2 The general case
More generally, the state $\hat{\rho}$ may have components with non-maximal spin length $S$. In this case, the state $\hat{\rho}$ can no longer be faithfully represented by a probability distribution on a sphere. Nonetheless, we meaningfully visualize the spin polarization of $\hat{\rho}$ by averaging over probability distributions that represent components of $\hat{\rho}$ within fixed-$S$ manifolds. To this end, we classify states by their spin length $S$, spin projection $m$ onto a quantization axis, and an auxiliary index $\xi$ that encodes how a state transforms under spin permutations. That is, we identify states |$\ket{S,m,\xi}$ that are simultaneous eigenstates of $\hat{S}^{2}=\hat{S}_{x}^{2}+\hat{S}_{y}^{2}+\hat{S}_{z}^{2}$ and $\hat{S}_{z}$, with $\hat{S}^{2}\ket{S,m,\xi}=S(S+1)\ket{S,m,\xi}$ and $\hat{S}_{z}\ket{S,m,\xi}=m\ket{S,m,\xi}$. We then define the rotated state
|$\ket{\bm{v},S,\xi}=\hat{R}(\bm{v})\ket{S,S,\xi},$ (15)
which is the analogue of |$\ket{\bm{v}}$ in a sector of Hilbert space with spin length $S$. The net spin-polarization probability distribution $\rho(\bm{v})$ is then
$\rho(\bm{v})=\sum_{S,\xi}\,w_{S}\braket{\bm{v},S,\xi}{\hat{\rho}}{\bm{v},S,\xi}.$ (16)